# BlurDM: A Blur Diffusion Model for Image Deblurring

### v2

@@ Abstract
Diffusion models show promise for dynamic scene deblurring; however, existing studies often fail to leverage the intrinsic nature of the blurring process within diffusion models, limiting their full potential. To address it, we present a Blur Diffusion Model (BlurDM), which seamlessly integrates the blur formation process into diffusion for image deblurring. Observing that motion blur stems from continuous exposure, BlurDM implicitly models the blur formation process through a dual-diffusion forward scheme, diffusing both noise and blur onto a sharp image. During the reverse generation process, we derive a dual denoising and deblurring formulation, enabling BlurDM to recover the sharp image by simultaneously denoising and deblurring, given pure Gaussian noise conditioned on the blurred image as input. Additionally, to efficiently integrate BlurDM into deblurring networks, we perform BlurDM in the latent space, forming a flexible prior generation network for deblurring. Extensive experiments demonstrate that BlurDM significantly and consistently enhances existing deblurring methods on four benchmark datasets. The project page is available at https://jin-ting-he.github.io/Blur-Diffusion-Model/.
@@ 1 Introduction
Camera shake or moving objects frequently introduce unwanted blur artifacts in captured images, severely degrading image quality and hindering downstream vision applications, such as object detection Kim_2024_CVPR; Wang_2024_CVPR, semantic segmentation Benigmim_2024_CVPR; Weber_2024_CVPR, and face recognition Minchul_2024_CVPR; Mi_2024_CVPR. Dynamic scene image deblurring aims to restore sharp details from a single blurred image, a highly ill-posed problem due to the directional and non-uniform nature of blur.
With the advancement of deep learning, CNN-based models gao2019dynamic; Nah_2017_CVPR; MT_2020_ECCV; tao2018srndeblur; Zamir2021MPRNet; Zhang_2019_CVPR have demonstrated remarkable success in data-driven deblurring. Additionally, Transformer-based approaches IPT; Kong_2023_CVPR; mao2024loformer; Tsai2022Stripformer; Wang_2022_CVPR; Zamir2021Restormer have been introduced to effectively capture long-range dependencies, further enhancing deblurring performance by leveraging global contextual information. Although previous methods have successfully improved deblurring performance, the inherent constraints of regression loss chen2023hierarchical typically lead to over-smoothed results with limited high-frequency details.
Recent advances in diffusion models NEURIPS2020_4c5bcfec; Rombach_2022_CVPR; song2020denoising have demonstrated remarkable success in image generation, producing high-quality images with rich details and sharp textures through a forward noise diffusion followed by reverse denoising. Building on the success of diffusion models, several studies chen2023hierarchical; Liu_2024_CVPR; rao2024rethinking; Ren_2023_ICCV; xia2023diffir have incorporated them into deblurring models to produce restored images. However, standard diffusion models are not specifically designed for deblurring. Thus, directly applying them to deblurring networks limits their potential, leading to suboptimal performance.
@@
In the reverse generation process, BlurDM aims to simultaneously remove noise and blur to restore sharp images. To overcome the challenge of unavailable ground-truth blur residuals, we derive a dual denoising and deblurring formulation that follows the principles of the blur formation process to implicitly approximate noise and blur residuals through dedicated noise and blur residual estimators. By effectively reversing the blur formation process, BlurDM can generate high-quality, realistic, sharp images. However, when applied to image deblurring, diffusion models may struggle to accurately reconstruct details with high content fidelity due to their inherent stochastic nature ye2024learning. To address the limitations, inspired by chen2023hierarchical; xia2023diffir, we utilize BlurDM as a prior generation network to flexibly and efficiently enhance existing deblurring models. Guided by the priors learned via BlurDM, deblurred models can achieve more accurate and visually consistent results. Key contributions of this work are summarized as follows:
@@ 2.1 Image Deblurring
Image deblurring has made substantial progress with the development of deep learning. Numerous studies have explored CNN-based deblurring using recurrent architectures, such as multi-scale Nah_2017_CVPR; tao2018srndeblur, multi-patch zamir2021multi; Zhang_2019_CVPR, and multi-temporal MT_2020_ECCV recurrent networks. For example, Tau et al.  tao2018srndeblur develop a scale-recurrent network accompanied by a coarse-to-fine strategy for deblurring. Zamir et al.  zamir2021multi introduce a multi-stage patch-recurrent network that splits an image into non-overlapping patches for hierarchical blurred pattern handling. Park et al. MT_2020_ECCV designs a temporal-recurrent network that progressively recovers sharp images through incremental temporal training.
Transformer-based methods IPT; Kong_2023_CVPR; mao2024loformer; Tsai2022Stripformer; Wang_2022_CVPR; Zamir2021Restormer have recently garnered considerable attention for deblurring due to their capacities to model long-range dependencies. However, the substantial training data and memory requirements of Transformers motivate the development of efficient variants Kong_2023_CVPR; mao2024loformer; Tsai2022Stripformer; Wang_2022_CVPR; Zamir2021Restormer specifically tailored for deblurring. For instance, Zamir et al. Zamir2021Restormer introduced a channel-wise attention mechanism to reduce memory overhead. Tsai et al. Tsai2022Stripformer proposed a strip-wise attention mechanism to handle blurred patterns with diverse orientations and magnitudes. Kong et al. Kong_2023_CVPR presented frequency attention to replace dot product operations in the spatial domain with element-wise multiplications in the frequency domain. Mao et al. mao2024loformer incorporated local channel-wise attention in the frequency domain to capture cross-covariance in the attention mechanism.
Although the aforementioned advances have improved deblurring performance through various architectural and algorithmic designs, the inherent constraints of using the regression loss chen2023hierarchical frequently lead to over-smoothed results with limited high-frequency details, producing suboptimal deblurred images.
@@ 2.2 Diffusion Models
Diffusion models NEURIPS2020_4c5bcfec; song2021denoising have demonstrated remarkable capability in generating high-fidelity images with rich details through forward noise diffusion and reverse denoising. They have been leveraged in numerous studies ma2024subject; mengsdedit; nichol2022glide; ramesh2021zero; ruiz2023dreambooth; saharia2022photorealistic; 5540175; ID-Blau; zhang2023adding to synthesize high-quality images under a variety of conditioning schemes.
Diffusion models have been applied to low-level vision tasks garber2024image; he2024domain; li2024rethinking; liu2024residual; liu2024diff; xia2023diffir; Zheng_2024_CVPR. For instance, Xia et al. xia2023diffir employed diffusion models to generate prior representations for clean image recovery. Liu et al. liu2024diff utilized text prompts to compile task-specific priors across various image restoration tasks. Zheng et al. Zheng_2024_CVPR proposed a selective hourglass mapping strategy to learn shared information between different tasks for universal image restoration.
Recognizing the advances of diffusion models in low-level vision, researchers have extended their use to image deblurring chen2024efficient; chen2023hierarchical; li2024hierarchical; liang2025swin; Liu_2024_CVPR; rao2024rethinking; ren2023multiscale; whang2022deblurring. Specifically, Whang et al. whang2022deblurring introduced a stochastic refinement diffusion model for deblurring. Ren et al. ren2023multiscale incorporated a multi-scale structure guidance network within the diffusion model to recover sharp images. Furthermore, several studies chen2024efficient; chen2023hierarchical; liang2025swin; rao2024rethinking employed diffusion models as prior generation networks and perform diffusion in the latent space to improve deblurring efficiency. For instance, Chen et al. chen2023hierarchical proposed a hierarchical integration module to fuse diffusion priors for deblurring, while Chen et al. chen2024efficient incorporated these priors into window-based transformer blocks. While these methods effectively reduce diffusion model latency for deblurring, they overlook the intrinsic characteristics of the blurring process within the diffusion framework, limiting their full potential.
Although Liu et al. Liu_2024_CVPR proposed residual diffusion by computing the difference between sharp and blurred images using a subtraction operation, the blur formation process is inherently a convolutional process rather than a direct additive difference, making this approach insufficient for accurately capturing blur characteristics. In contrast, we propose a novel framework that incorporates the blur formation process into the diffusion model, leading to significant deblurring performance improvements.
@@ 3 Proposed Method
We propose Blur Diffusion Model (BlurDM), a novel diffusion framework for image deblurring. Unlike existing methods chen2023hierarchical; rao2024rethinking; Ren_2023_ICCV; xia2023diffir, which rely only on noise diffusion, BlurDM integrates a blur diffusion process, incorporating blur formation into diffusion to improve deblurring performance.
@@ 3.1 Dual Noise and Blur Diffusion Process
However, existing deblur datasets typically consist of blurry-sharp image pairs without providing the corresponding blur residuals. To address this limitation, we reparameterize (5) to obtain the conditional probability distribution $q(I_{T}|I_{0},e_{1:T})$ NEURIPS2020_4c5bcfec, as shown in (6). The full derivation is provided in Appendix A.1.
@@ 3.2 Dual Denoising and Deblurring Process
Inspired by the deterministic sampling formulation in DDIM song2021denoising, we define the reverse transition distribution as
@@ 3.3 Latent BlurDM
To efficiently integrate BlurDM into deblurring networks, we develop it in the latent space, where it serves as a prior generator to enhance existing deblurring methods. Inspired by previous work chen2023hierarchical; xia2023diffir, we adopt a three-stage training strategy for effective integration, as illustrated in Fig. 2. This process guides the latent features to capture physically meaningful representations for blur residuals, allowing the model to encode exposure-aware information in the latent space.
@@
which is used to estimate the sharp prior $Z^{S}$. Recent studies chen2023hierarchical; geng2025consistency; salimansprogressive; xia2023diffir reveal that supervision on the final output can effectively influence the entire diffusion trajectory. We define a latent‑prior loss $\mathcal{L}_{\text{prior}}=\left\lVert Z^{B}_{0}-Z^{S}\right\rVert_{1},$ where $Z^{B}_{0}$ is obtained by recursively removing estimated blur and noise residuals from $I_{T}$ via the shared estimators $e^{\theta}$ and $\epsilon^{\theta}$. By back‑propagating $\mathcal{L}_{\text{prior}}$, gradients are distributed across all reverse steps, furnishing amortised trajectory‑level supervision without step‑wise labels. Further details are provided in Appendix A.2.
@@ 4.1 Experimental Setup
We adopt four prominent deblurring models, including MIMO-UNet MIMO, Stripformer Tsai2022Stripformer, FFTformer Kong_2023_CVPR, and LoFormer mao2024loformer, to validate the effectiveness of BlurDM. Following previous work MIMO; Kong_2023_CVPR; mao2024loformer; Tsai2022Stripformer, we adopt the widely used GoPro Nah_2017_CVPR and HIDE HAdeblur datasets. The GoPro dataset contains $2,103$ image pairs for training and $1,111$ image pairs for testing, while the HIDE dataset contains $2,025$ image pairs used only for testing. Additionally, we utilize the real-world RealBlur rim_2020_ECCV dataset, which contains RealBlur-J and RealBlur-R subsets. Each subset contains $3,758$ training pairs and $980$ testing pairs, with RealBlur-J in JPEG and RealBlur-R in Raw format.
@@ 4.3 Ablation studies
We evaluate the performance of the baseline deblurring model enhanced by various prior generation methods, including MLP, DDPM NEURIPS2020_4c5bcfec, RDDM Liu_2024_CVPR, and our proposed BlurDM, on the GoPro and RealBlur-J datasets (see Tab. 3). “Net1” denotes the baseline model without guidance by a prior generation network. “Net2” denotes the deblurring model enhanced with MLP layers, without using a diffusion process. “Net3”, “Net4”, and “Net5” correspond to the deblurring models enhanced by different diffusion-based priors, including DDPM NEURIPS2020_4c5bcfec, RDDM Liu_2024_CVPR, and the proposed BlurDM, respectively.
@@
Compared to the standard diffusion model in NEURIPS2020_4c5bcfec, which requires thousands of iterations in the reverse generation process, applying diffusion networks in the latent space has proven effective in reducing the number of iterations chen2023hierarchical; xia2023diffir. Therefore, we examine the effect of different iteration counts used in BlurDM on deblurring performance, as shown in Fig. 7. Specifically, we test eight iteration settings $T\in\{0,1,2,4,5,6,8,10\}$ and evaluate their deblurring performances on the GoPro test set. The results show that BlurDM significantly improves performance with two iterations and reaches peak performance after five, showcasing its ability to achieve substantial and stable performance gains with only a few iterations.
@@
We compare BlurDM with two recent diffusion-based approaches, HI-Diff chen2023hierarchical and RDDM liu2024residual, in Tab. 6. HI-Diff follows the conventional diffusion process in the latent space, where Gaussian noise is progressively injected directly into the latent until it becomes pure noise, and a learned reverse process is used to reconstruct a clean latent for deblurring. RDDM first forms an image space residual by subtracting the clean image from the degraded one, then performs diffusion on the clean image that jointly models this residual and Gaussian noise. Both HI-Diff and RDDM neglect the physics of blur formation. BlurDM addresses this gap by explicitly integrating the blur formation process with diffusion, executing dual noise and blur diffusion that matches the physics of blur accumulation. With comparable parameter counts and FLOPs, BlurDM consistently delivers higher PSNR and SSIM. As a plug-and-play module, it integrates seamlessly with diverse backbone architectures, demonstrating both strong performance and broad generalizability.

### v3

@@ Abstract
Diffusion models show promise for dynamic scene deblurring; however, existing studies often fail to leverage the intrinsic nature of the blurring process within diffusion models, limiting their full potential. To address it, we present a Blur Diffusion Model (BlurDM), which seamlessly integrates the blur formation process into diffusion for image deblurring. Observing that motion blur stems from continuous exposure, BlurDM implicitly models the blur formation process through a dual-diffusion forward scheme, diffusing both noise and blur onto a sharp image. During the reverse generation process, we derive a dual denoising and deblurring formulation, enabling BlurDM to recover the sharp image by simultaneously denoising and deblurring, given pure Gaussian noise conditioned on the blurred image as input. Additionally, to efficiently integrate BlurDM into deblurring networks, we perform BlurDM in the latent space, forming a flexible prior generation network for deblurring. Extensive experiments demonstrate that BlurDM significantly and consistently enhances existing deblurring methods on four benchmark datasets. The project page is available at https://jin-ting-he.github.io/BlurDM/.
@@ 1 Introduction
Camera shake or moving objects frequently introduce unwanted blur artifacts in captured images, severely degrading image quality and hindering downstream vision applications, such as object detection Kim et al. (2024a); Wang et al. (2024), semantic segmentation Benigmim et al. (2024); Weber et al. (2024), and face recognition Kim et al. (2024b); Mi et al. (2024). Dynamic scene image deblurring aims to restore sharp details from a single blurred image, a highly ill-posed problem due to the directional and non-uniform nature of blur.
With the advancement of deep learning, CNN-based models Gao et al. (2019); Nah et al. (2017); Park et al. (2020); Tao et al. (2018); Zamir et al. (2021a); Zhang et al. (2019) have demonstrated remarkable success in data-driven deblurring. Additionally, Transformer-based approaches Chen et al. (2021); Kong et al. (2023); Mao et al. (2024); Tsai et al. (2022); Wang et al. (2022); Zamir et al. (2022) have been introduced to effectively capture long-range dependencies, further enhancing deblurring performance by leveraging global contextual information. Although previous methods have successfully improved deblurring performance, the inherent constraints of regression loss Chen et al. (2023) typically lead to over-smoothed results with limited high-frequency details.
Recent advances in diffusion models Ho et al. (2020); Rombach et al. (2022); Song et al. (2021a) have demonstrated remarkable success in image generation, producing high-quality images with rich details and sharp textures through a forward noise diffusion followed by reverse denoising. Building on the success of diffusion models, several studies Chen et al. (2023); Liu et al. (2024a); Rao et al. (2024); Ren et al. (2023a); Xia et al. (2023) have incorporated them into deblurring models to produce restored images. However, standard diffusion models are not specifically designed for deblurring. Thus, directly applying them to deblurring networks limits their potential, leading to suboptimal performance.
@@
In the reverse generation process, BlurDM aims to simultaneously remove noise and blur to restore sharp images. To overcome the challenge of unavailable ground-truth blur residuals, we derive a dual denoising and deblurring formulation that follows the principles of the blur formation process to implicitly approximate noise and blur residuals through dedicated noise and blur residual estimators. By effectively reversing the blur formation process, BlurDM can generate high-quality, realistic, sharp images. However, when applied to image deblurring, diffusion models may struggle to accurately reconstruct details with high content fidelity due to their inherent stochastic nature Ye et al. (2024). To address the limitations, inspired by Chen et al. (2023); Xia et al. (2023), we utilize BlurDM as a prior generation network to flexibly and efficiently enhance existing deblurring models. Guided by the priors learned via BlurDM, deblurred models can achieve more accurate and visually consistent results. Key contributions of this work are summarized as follows:
@@ 2.1 Image Deblurring
Image deblurring has made substantial progress with the development of deep learning. Numerous studies have explored CNN-based deblurring using recurrent architectures, such as multi-scale Nah et al. (2017); Tao et al. (2018), multi-patch Zamir et al. (2021b); Zhang et al. (2019), and multi-temporal Park et al. (2020) recurrent networks. For example, Tau et al.  Tao et al. (2018) develop a scale-recurrent network accompanied by a coarse-to-fine strategy for deblurring. Zamir et al.  Zamir et al. (2021b) introduce a multi-stage patch-recurrent network that splits an image into non-overlapping patches for hierarchical blurred pattern handling. Park et al. Park et al. (2020) designs a temporal-recurrent network that progressively recovers sharp images through incremental temporal training.
Transformer-based methods Chen et al. (2021); Kong et al. (2023); Mao et al. (2024); Tsai et al. (2022); Wang et al. (2022); Zamir et al. (2022) have recently garnered considerable attention for deblurring due to their capacities to model long-range dependencies. However, the substantial training data and memory requirements of Transformers motivate the development of efficient variants Kong et al. (2023); Mao et al. (2024); Tsai et al. (2022); Wang et al. (2022); Zamir et al. (2022) specifically tailored for deblurring. For instance, Zamir et al. Zamir et al. (2022) introduced a channel-wise attention mechanism to reduce memory overhead. Tsai et al. Tsai et al. (2022) proposed a strip-wise attention mechanism to handle blurred patterns with diverse orientations and magnitudes. Kong et al. Kong et al. (2023) presented frequency attention to replace dot product operations in the spatial domain with element-wise multiplications in the frequency domain. Mao et al. Mao et al. (2024) incorporated local channel-wise attention in the frequency domain to capture cross-covariance in the attention mechanism.
Although the aforementioned advances have improved deblurring performance through various architectural and algorithmic designs, the inherent constraints of using the regression loss Chen et al. (2023) frequently lead to over-smoothed results with limited high-frequency details, producing suboptimal deblurred images.
@@ 2.2 Diffusion Models
Diffusion models Ho et al. (2020); Song et al. (2021b) have demonstrated remarkable capability in generating high-fidelity images with rich details through forward noise diffusion and reverse denoising. They have been leveraged in numerous studies Ma et al. (2024); Meng et al. (2022); Nichol et al. (2022); Ramesh et al. (2021); Ruiz et al. (2023); Saharia et al. (2022); Whyte et al. (2010); Wu et al. (2024); Zhang et al. (2023) to synthesize high-quality images under a variety of conditioning schemes.
Diffusion models have been applied to low-level vision tasks Garber and Tirer (2024); He et al. (2024); Li et al. (2024a); Liu et al. (2024b, c); Xia et al. (2023); Zheng et al. (2024). For instance, Xia et al. Xia et al. (2023) employed diffusion models to generate prior representations for clean image recovery. Liu et al. Liu et al. (2024c) utilized text prompts to compile task-specific priors across various image restoration tasks. Zheng et al. Zheng et al. (2024) proposed a selective hourglass mapping strategy to learn shared information between different tasks for universal image restoration.
Recognizing the advances of diffusion models in low-level vision, researchers have extended their use to image deblurring Chen and Liu (2024); Chen et al. (2023); Li et al. (2024b); Liang et al. (2025); Liu et al. (2024a); Rao et al. (2024); Ren et al. (2023b); Whang et al. (2022). Specifically, Whang et al. Whang et al. (2022) introduced a stochastic refinement diffusion model for deblurring. Ren et al. Ren et al. (2023b) incorporated a multi-scale structure guidance network within the diffusion model to recover sharp images. Furthermore, several studies Chen and Liu (2024); Chen et al. (2023); Liang et al. (2025); Rao et al. (2024) employed diffusion models as prior generation networks and perform diffusion in the latent space to improve deblurring efficiency. For instance, Chen et al. Chen et al. (2023) proposed a hierarchical integration module to fuse diffusion priors for deblurring, while Chen et al. Chen and Liu (2024) incorporated these priors into window-based transformer blocks. While these methods effectively reduce diffusion model latency for deblurring, they overlook the intrinsic characteristics of the blurring process within the diffusion framework, limiting their full potential.
Although Liu et al. Liu et al. (2024a) proposed residual diffusion by computing the difference between sharp and blurred images using a subtraction operation, the blur formation process is inherently a convolutional process rather than a direct additive difference, making this approach insufficient for accurately capturing blur characteristics. In contrast, we propose a novel framework that incorporates the blur formation process into the diffusion model, leading to significant deblurring performance improvements.
@@ 3 Proposed Method
We propose Blur Diffusion Model (BlurDM), a novel diffusion framework for image deblurring. Unlike existing methods Chen et al. (2023); Rao et al. (2024); Ren et al. (2023a); Xia et al. (2023), which rely only on noise diffusion, BlurDM integrates a blur diffusion process, incorporating blur formation into diffusion to improve deblurring performance.
@@ 3.1 Dual Noise and Blur Diffusion Process
However, existing deblur datasets typically consist of blurry-sharp image pairs without providing the corresponding blur residuals. To address this limitation, we reparameterize (5) to obtain the conditional probability distribution $q(I_{T}|I_{0},e_{1:T})$ Ho et al. (2020), as shown in (6). The full derivation is provided in Appendix A.1.
@@ 3.2 Dual Denoising and Deblurring Process
Inspired by the deterministic sampling formulation in DDIM Song et al. (2021b), we define the reverse transition distribution as
@@ 3.3 Latent BlurDM
To efficiently integrate BlurDM into deblurring networks, we develop it in the latent space, where it serves as a prior generator to enhance existing deblurring methods. Inspired by previous work Chen et al. (2023); Xia et al. (2023), we adopt a three-stage training strategy for effective integration, as illustrated in Fig. 2. This process guides the latent features to capture physically meaningful representations for blur residuals, allowing the model to encode exposure-aware information in the latent space.
@@
which is used to estimate the sharp prior $Z^{S}$. Recent studies Chen et al. (2023); Geng et al. (2025); Salimans and Ho (2022); Xia et al. (2023) reveal that supervision on the final output can effectively influence the entire diffusion trajectory. We define a latent‑prior loss $\mathcal{L}_{\text{prior}}=\left\lVert Z^{B}_{0}-Z^{S}\right\rVert_{1},$ where $Z^{B}_{0}$ is obtained by recursively removing estimated blur and noise residuals from $I_{T}$ via the shared estimators $e^{\theta}$ and $\epsilon^{\theta}$. By back‑propagating $\mathcal{L}_{\text{prior}}$, gradients are distributed across all reverse steps, furnishing amortised trajectory‑level supervision without step‑wise labels. Further details are provided in Appendix A.2.
@@ 4.1 Experimental Setup
We adopt four prominent deblurring models, including MIMO-UNet Cho et al. (2021), Stripformer Tsai et al. (2022), FFTformer Kong et al. (2023), and LoFormer Mao et al. (2024), to validate the effectiveness of BlurDM. Following previous work Cho et al. (2021); Kong et al. (2023); Mao et al. (2024); Tsai et al. (2022), we adopt the widely used GoPro Nah et al. (2017) and HIDE Shen et al. (2019) datasets. The GoPro dataset contains $2,103$ image pairs for training and $1,111$ image pairs for testing, while the HIDE dataset contains $2,025$ image pairs used only for testing. Additionally, we utilize the real-world RealBlur Rim et al. (2020) dataset, which contains RealBlur-J and RealBlur-R subsets. Each subset contains $3,758$ training pairs and $980$ testing pairs, with RealBlur-J in JPEG and RealBlur-R in Raw format.
@@ 4.3 Ablation studies
We evaluate the performance of the baseline deblurring model enhanced by various prior generation methods, including MLP, DDPM Ho et al. (2020), RDDM Liu et al. (2024a), and our proposed BlurDM, on the GoPro and RealBlur-J datasets (see Tab. 3). “Net1” denotes the baseline model without guidance by a prior generation network. “Net2” denotes the deblurring model enhanced with MLP layers, without using a diffusion process. “Net3”, “Net4”, and “Net5” correspond to the deblurring models enhanced by different diffusion-based priors, including DDPM Ho et al. (2020), RDDM Liu et al. (2024a), and the proposed BlurDM, respectively.
@@
Compared to the standard diffusion model in Ho et al. (2020), which requires thousands of iterations in the reverse generation process, applying diffusion networks in the latent space has proven effective in reducing the number of iterations Chen et al. (2023); Xia et al. (2023). Therefore, we examine the effect of different iteration counts used in BlurDM on deblurring performance, as shown in Fig. 7. Specifically, we test eight iteration settings $T\in\{0,1,2,4,5,6,8,10\}$ and evaluate their deblurring performances on the GoPro test set. The results show that BlurDM significantly improves performance with two iterations and reaches peak performance after five, showcasing its ability to achieve substantial and stable performance gains with only a few iterations.
@@
We compare BlurDM with two recent diffusion-based approaches, HI-Diff Chen et al. (2023) and RDDM Liu et al. (2024b), in Tab. 6. HI-Diff follows the conventional diffusion process in the latent space, where Gaussian noise is progressively injected directly into the latent until it becomes pure noise, and a learned reverse process is used to reconstruct a clean latent for deblurring. RDDM first forms an image space residual by subtracting the clean image from the degraded one, then performs diffusion on the clean image that jointly models this residual and Gaussian noise. Both HI-Diff and RDDM neglect the physics of blur formation. BlurDM addresses this gap by explicitly integrating the blur formation process with diffusion, executing dual noise and blur diffusion that matches the physics of blur accumulation. With comparable parameter counts and FLOPs, BlurDM consistently delivers higher PSNR and SSIM. As a plug-and-play module, it integrates seamlessly with diverse backbone architectures, demonstrating both strong performance and broad generalizability.